OpenStudy (anonymous):

Solve the quadratic equation below and describe what the solution represents on a graph Show your work so i can understand 0 = x2 + 5x + 6

OpenStudy (mathteacher1729):

Megan, are you allowed to use a graphing calc on this problem?

OpenStudy (anonymous):

no sadly not

OpenStudy (anonymous):

i just need to know the points on the vertex

OpenStudy (mathteacher1729):

Ok. Well, you can factor this. :) Do you know how to factor a quadratic?

OpenStudy (anonymous):

yes but i dont understand the factors of 5 and 6

OpenStudy (mathteacher1729):

Well you want to write the equation like this: \[(x +a )(x +b)\] Where a times b = 6 and a plus b = 5 So the question is ... what two numbers make that true?

OpenStudy (anonymous):

oh sorry 3 and 2 multiply to get 6 and add up to 5 i understand a lil better now

OpenStudy (mathteacher1729):

Yay! Ok, so we're almost there. :) \[(x+3)(x+2) = x^2+5x+6\] So when is \((x+3)(x+2)=0)\) ? Hint it's wherever (x+3) = 0 or (x+2) = 0

OpenStudy (anonymous):

x= -3? and x= -2?

OpenStudy (mathteacher1729):

Yup! So now we know that this quadratic equation intersects the x axis at x = -2 and x = -3. Since the \(x^2\) term is positive, that means the parabola opens upward. The vertex is at the x value EXACTLY between -2 and -3.

OpenStudy (anonymous):

so is that the answer?

OpenStudy (mathteacher1729):

Well, you solved the qudratic. (found the values of x which made it true). As for the graph, can you now sketch the parbola knowing what we found out?

OpenStudy (anonymous):

x^2 + 5x + 6=0 x^2+(3x+2x) +6=0 (x^2+3x)+(2x+6)=0 x(x+3)+2(x+3)=0 (x+3)(x+2)=0 either , (x+3)=0 or, (x+2)=0 therefore, x =-3 or x =-2

OpenStudy (anonymous):

yes i think so do i need y?

OpenStudy (mathteacher1729):

Well if you're going to sketch it, you need an x an y axis. :)

OpenStudy (anonymous):

if i want to get the y dont i just take one of the x's and put it in the equation?

OpenStudy (mathteacher1729):

Yes. So when x = - 3 or - 2, y will be zero. What about when x = -2.5 ? What will y equal?

OpenStudy (anonymous):

y=31?

OpenStudy (phi):

This is cheating a bit, but it gives a clue. http://www.wolframalpha.com/input/?i=y+%3D+x%5E2+%2B+5x+%2B+6+

OpenStudy (anonymous):

when x = -2.5 , then y=0

OpenStudy (anonymous):

oh so y just equals 0?

OpenStudy (mathteacher1729):

"when x = -2.5 , then y=0 " This is incorrect. y only equals 0 for x = -2 and x = -3. For all other values of x, y is nonzero.

OpenStudy (mathteacher1729):

To see what this graph looks like, go to http://www.wolframalpha.com and type in x^2 + 5x + 6 for x = - 4 to -1

OpenStudy (anonymous):

so on the graph i jus put (-2.5,0)

OpenStudy (phi):

Unfortunately, when x= -2.5 (or -5/2) y does not equal 0 Try again!

OpenStudy (anonymous):

so could i just but like (-2,0)?

OpenStudy (phi):

I would plot (-2,0). (-3, 0) and the middle point at (-5/2 , ??) and sketch a U shaped curve. Try to find y when x= -5/2.

OpenStudy (anonymous):

1/2((-2,0)+(-3, 0))= ((-2-3)/2,(0+0)/2) =(-5/2,0)

OpenStudy (phi):

@ab we are trying to plot y=x^2 + 5x + 6.

OpenStudy (anonymous):

ok i think i will plot it at (-2,0) and (-3,0)

OpenStudy (phi):

y= (-5/2)^2 + 5(-5/2) +6 y= 25/4 - 25/2 +6 Can you finish?

OpenStudy (phi):

Another easy point to find is when x=0 y= 0^2 +5*0 +6 =6 so (0,6) is also on the curve

OpenStudy (anonymous):

ok thanks yah i can finishh